# Immersed figure-8 annuli and anyons

### v2

@@ 5.2 Constraints of transportation
As a corollary, $|{\mathcal{C}}|=|{\mathcal{C}}_{8}|$. This is because transporting an anyon along a figure-8 immersed in a ball or sphere cannot change the anyon type.
Here is a sketch of the key idea of proof of Prop. 9. The trick is to attach the test annulus to another annulus by a strip. After transporting the test annulus, the strip will be stretched along loop ${\mathcal{L}}$. After gluing the test annulus with the other annulus (by merging with another annulus), we get a punctured torus ${\mathcal{W}}_{\mathcal{L}}$. This process is illustrated in Fig. 11. The main part of the proof is to analyze $\Sigma({\mathcal{W}}_{\mathcal{L}})$, or more conveniently $\Sigma(T^{2})$, where the torus $T^{2}$ has a reference state obtained by an analog of “vacuum block completion” (see Example 4.27 of [17]).
@@
Consider the process illustrated in Fig. 11. We start with a disk away from the defect points (black dots) and remove two small balls from it. Let the resulting 2-hole disk be $Y$. Consider a state labeled by $a,1,a$ on the three boundaries $\rho^{a1a}_{Y}\in\Sigma(Y)$. It must be an extreme point because the corresponding fusion space is 1 dimensional.
@@
Figure 11: Construction of an extreme point in $\Sigma_{\hat{1}}({\mathcal{W}}_{\mathcal{L}})$, useful in the proof of Prop. 9.
If $a\in{\mathcal{I}}({\mathcal{L}})$, that is, $a$ is invariant under the transportation along ${\mathcal{L}}$, we do step 3 of Fig. 11, which glues annulus $X$ with the state on $Y_{\mathcal{L}}$ to a punctured torus ${\mathcal{W}}_{\mathcal{L}}$. In terms of quantum states, we merge two quantum states, which are extreme points of $\Sigma(X)$ and $\Sigma(Y_{\mathcal{L}})$ (thickening a certain region before merging if needed) with the superselection sectors labeled in Fig. 11(c). Here, $a\in{\mathcal{I}}({\mathcal{L}})$ is required so that the two states match on their overlap regions. For each valid choice of $a$, we get an extreme point of $\Sigma_{\hat{1}}({\mathcal{W}}_{\mathcal{L}})$. Note that $1\in{\mathcal{I}}({\mathcal{L}})$, and therefore $\Sigma_{\hat{1}}({\mathcal{W}}_{\mathcal{L}})$ is nonempty. Note that the punctured torus ${\mathcal{W}}_{\mathcal{L}}$ contains annuli $X_{\mathcal{L}}$ and an embedded test annulus $X$.
@@
where $\delta_{a,b}$ comes from the matching of sector on one shared boundary of the two annuli, and $\delta_{a,\varphi_{\mathcal{L}}(b)}$ comes from the other shared boundary. The appearance of $\varphi_{\mathcal{L}}$ is due to the transportation property.
Similarly, we cut the torus along another direction into two annuli, $T^{2}=X_{\mathcal{L}}\bar{X}_{\mathcal{L}}$. Here, $X_{\mathcal{L}}$ is the immersed annulus that thickens ${\mathcal{L}}$, a notation explained in Fig. 9. Not only are $X_{\mathcal{L}}$ and $\bar{X}_{\mathcal{L}}$ both sectorizable, but they must also have isomorphism information convex sets $\Sigma(X_{\mathcal{L}})\cong\Sigma(\bar{X}_{\mathcal{L}})$. This is because $X_{\mathcal{L}}$ deforms to $\bar{X}_{\mathcal{L}}$ smoothly on $T^{2}$. So they both characterize ${\mathcal{C}}_{\mathcal{L}}$. From a computation parallel to Eq. (28), we see
@@
An annulus is a disk with a puncture, homeomorphically speaking. If the state on $X_{\mathcal{L}}$ is an extreme point, we can use the completion trick (Lemma 4.4 of [17]) to heal a puncture to obtain a state on the disk. The resulting state satisfies the axioms on the entire disk if the extreme point is associated with an Abelian sector ($\mu\in{\mathcal{C}}_{\mathcal{L}}$ with $d_{\mu}=1$). Let this state be $\tilde{\sigma}$. It is known that transporting a test annulus on a disk cannot permute anyon types, supposing that the disk has a reference state. (This is by Lemma 4.3 of [14]). As the loop $X_{\mathcal{L}}$ is identified as part of the disk for which we have a valid reference state $\tilde{\sigma}$, the transportation of test annulus on $X_{\mathcal{L}}$ cannot permute anyons. Thus, the transportation along ${\mathcal{L}}$ is parallel. ∎
@@ 6 Discussion
In this work, we started by asking a simple question: “is there an Abelian superselection sector detected by the immersed figure-8 annulus?” The question is relevant to gapped topologically ordered systems in 2D, which support anyons and is formulated in the framework of entanglement bootstrap. We gave an affirmative answer to this question for Abelian anyon theories but leave the general case as a conjecture (Conjecture 5). We explained the fact that if there is an Abelian sector on figure-8, a certain strong isomorphism (Conjecture 6) of information convex sets must be true: if two homeomorphic regions $\Omega$ and $\Omega^{\prime}$ are immersed in the sphere or a disk, the information convex sets $\Sigma(\Omega)$ and $\Sigma(\Omega^{\prime})$ must be isomorphic. Importantly, it does not matter if $\Omega$ can be smoothly deformed to $\Omega^{\prime}$ on the sphere (or disk), a situation where the established isomorphism theorem does not apply.
@@
Suppose the existence of an Abelian sector on figure-8 (Conjecture 5) can be proved. Strong isomorphism will be a fact. While strong isomorphism is powerful, it seems to imply that we cannot learn much by looking at a topological space immersed differently. Let me argue this is far from the case. If strong isomorphism holds, a couple of intriguing questions can be pursued. In particular, it means tunneling processes (Section 3.1 and Fig. 5) can generate explicit isomorphisms not allowed by smooth deformation. Thus, automorphisms of information convex sets will not only come from topologically nontrivial ways to smoothly deform an immersed region back to itself but also from ways that utilize tunneling in intermediate steps. We expect these diverse topological classes of automorphisms to be informative. A natural question is: “Can we use these automorphisms (dancing of quantum states) to extract all the data of the modular tensor category underlying an anyon model?” [^14] An optimism to this possibility was found independent by Kyle Kawagoe, who attended my chalk talk at math department Ohio State University. If the answer is affirmative, it will manifestly use a single quantum state, and this will be complementary to the “microscopic approach” [37]. Can the mathematical structure coming out of immersion be comparable with the mapping class group? Analogous questions in higher dimensions can be asked in parallel and are more open.
A stronger conjecture we postponed until now is the existence of a canonical (Abelian) vacuum state on the immersed figure-8 annulus. If this is true, the process in Fig. 15 of the appendix may well be on the right track in extracting the full set of anyon topological spins. This is a meaningful problem that deserves an effort elsewhere. Proving the existence of a canonical vacuum on figure-8 remains an open problem, even for Abelian anyon theories, in which context we proved Conjecture 5.
@@
Finally, we ask whether considering quantum states on regions immersed in a physical system can be useful in other physical contexts. Immersed regions utilize the Hilbert space more efficiently by using local pieces more than once, and they can be made larger and topologically (or geometrically) more interesting than the original physical system available to us. Smooth deformation of immersion region, a property natural in systems with entanglement area law, will be absent in broader contexts such as gapless systems. Nevertheless, many other benefits of immersion may persist.
@@ Appendix A Equivalent definitions of quantum dimension
Immersed sphere shells, in the context of $n$-dimensional entanglement bootstrap, are sectorizable regions. We derive an equivalence relation on the definition of quantum dimensions (Lemma 12) for immersed sphere shells in $n$-dimensions, where $n\geq 2$ and is arbitrary. It applies to immersed annuli as a special case ($n=2$).
Because the setup is general, we introduce some diagrammatic notation for the partitions. Let $Z=S^{n-1}\times\mathbb{I}$ be a sphere shell in $n$ space dimensions, where $S^{n-1}$ is a $(n-1)$-sphere and $\mathbb{I}$ is an interval. In Fig. 12, the horizontal direction indicates the direction of the interval $\mathbb{I}$. Therefore, each vertical strip that connects the top and the bottom represents a sphere shell. If a strip is cut in half (top and bottom), each piece is a half-sphere shell (topologically a ball). We denote the superselection sectors detected by $\Sigma(Z)$ as ${\mathcal{C}}_{Z}$.
@@
The following are equivalent definitions of quantum dimension $d_{h}$ of $h\in{\mathcal{C}}_{Z}$, where $Z$ is an immersed sphere shell in $n$-dimension.
@@
We shall see in the proof that we do not need to deform the annulus. Therefore, Lemma 12 holds not only for sphere shells immersed in a sphere but also applicable to contexts allowing topological defects (i.e., when the reference state is on a manifold with a nontrivial topology).
@@
The region $Z$, as a sectorizable region, satisfies $\Delta(\partial Z,Z\setminus\partial Z)_{\rho^{h}}=0$, the so-called “extreme point criterion”. It implies
@@ Appendix B Automorphism of information convex sets: tools and examples
Automorphism (i.e., self-isomorphism) of information convex sets plays an important role in entanglement bootstrap. The simplest application is the definition of antiparticles, but its implication is much broader. In this appendix, we talk about two things related to automorphisms of information convex sets.
@@ B.1 Math related to immersion and deformation classes
Here, $\coprod$ means disjoint union, and $\mathfrak{m}(\Omega,N)$, for any $\Omega\looparrowright N$, is the simplicial 2-complex such that
@@
One may ask if it is meaningful to add higher dimensional $k$-simplexes ($k\geq 3$) instead of stopping at $k=2$. The relevance to physics we can think of so far (as we explain shortly) depends on $\pi_{0}$ and $\pi_{1}$. Adding higher dimensional simplexes will not change them. Nonetheless, it remains an interesting question if higher homotopy (or homology) groups of $\mathfrak{M}$ can provide further information. [^16] We thank John McGreevy for this question.
@@ B.1.1 Physical relevance of immersion complex through the homotopy groups
It is easy to explain why the number of connected components $\pi_{0}$ of $\mathfrak{M}(\omega,N)$ is of physical relevance. Consider the entanglement bootstrap setup where $N$ is the manifold on which we define our reference state $\sigma_{N}$. The $n$-dimensional version of axioms A0 and A1 are satisfied everywhere on $N$. Then $\pi_{0}(\mathfrak{M}(\omega,N))$, i.e., the number of connected components of $\mathfrak{M}(\omega,N)$, is the number of inequivalent immersions of the topological space $\omega$ in $N$. For instance, $\pi_{0}(\mathfrak{M}(\textrm{annulus},S^{2}))$ is $\mathbb{Z}_{2}$ (the cyclic group of order 2). We discussed this in the main text without using this notation. In broader contexts, $\pi_{0}(\mathfrak{M})$ is the set of regular homotopy class of immersed regions, which has been a topic of study in math literature; see, e.g., [38, 39, 28]. Within each connected component, $\mathfrak{m}(\Omega,N)$, $\Omega\in\mathcal{R}(\omega,N)$, we can smoothly deform the region and the isomorphism theorem implies that such deformation preserves the information convex set. Configurations in two connected components are harder to relate, although tunneling (Section 3.1) provides one way to generate isomorphisms between such configurations (when a certain Abelian sector exists).
@@ B.1.2 Examples
where $X(j)$ for $j\geq 0$ represents an immersed annulus with turning number $j$. In particular, we can let $X(0)=X_{8}$ be the immersed “figure-8” and let $X(1)=X$ be the embedded annulus. We expect $\mathfrak{m}(X,B^{2})$ to be simply connected and $\pi_{1}(\mathfrak{m}(X,B^{2}))$ is trivial. [^17] If we deform $X$ back to itself, keeping the intermediate configurations embedded, the deformation must correspond to a loop in $\mathfrak{m}(X,B^{2})$ that can shrink to a point. The argument is precisely that used to prove Lemma 4.3 of [14]. However, for the deformations allowing immersed intermediate configurations, the argument does not apply. What is $\pi_{1}(\mathfrak{m}(X_{8},B^{2}))$? From the intuition around Eq. (6), we conjecture that $\pi_{1}(\mathfrak{m}(X_{8},B^{2}))=\mathbb{Z}_{2}$. If this is true, the definition of anti-sector on $X_{8}$, discussed near Eq. (6) is justified.
@@
where $X$ and $X_{8}$ are the only two classes of immersed annuli in the sphere (as discussed in Table 1). Thus $\mathfrak{M}(\textrm{annulus},S^{2})$ has two connected components. What are the fundamental groups $\pi_{1}$ of the connected components? $\pi_{1}(\mathfrak{m}(X,S^{2}))$ should contain $\mathbb{Z}_{2}$ as a quotient group. The related observation is that $X$ can turn inside out on the sphere and the fact that such process can turn an anyon $a\in{\mathcal{C}}$ into anti-anyon $\bar{a}\in{\mathcal{C}}$, which is, in general, a different sector. (For a related discussion, see Appendix G of [14].) We conjecture that $\pi_{1}(\mathfrak{m}(X_{8},S^{2}))$ is $\mathbb{Z}_{2}$.
@@
where $Z$, on the right-hand side, represents the embedded sphere shell in the 3-dimensional ball $B^{3}$. Eq. (43) follows from the fact that there is a unique way to immerse a sphere $S^{2}$ in $B^{3}$ (or $R^{3}$), up to regular homotopy; this is an influential early result by Smale [38], which is also the topic of video [29]. It seems plausible that $\pi_{1}(\mathfrak{m}(Z,B^{3}))=\mathbb{Z}_{2}$, but we do not know proofs. It is further interesting to consider
@@
where $G_{g}$ is an embedded genus-$g$ handlebody and $g$ is a positive integer. As is indicated in Eq. (44), $\mathfrak{M}$ is connected in this example. Genus-$g$ handlebodies detect the “graph excitations” studied in [16, 17], where for $g=1$ these graph excitations are flux loops. In general, $\pi_{1}(\mathfrak{m}(G_{g},B^{3}))$ can have pretty nontrivial structure even though $G_{g}$ are sectorizable regions.
Figure 14: A deformation process, which corresponds to an oriented loop in $\mathfrak{m}(X,X_{+})$. Here, $X$ is an annulus (thick black loop in the top figure) embedded in a wider annulus $X_{+}$ (gray).
@@
Note that, $\pi_{1}(\mathfrak{m}(X,X_{+}))$ can be nontrivial even if $\mathfrak{m}(X,B^{2})$ is trivial. Intuitively, in the process described in Fig. 14, a pair of twists are created locally and then transported around the annulus before they annihilate again. (The process in Fig. 14 is inspired by Fig. 8 of [39] as well as Fig. 7 of [20].) Can this process represent a nontrivial element of $\pi_{1}(\mathfrak{m}(X,X_{+}))$? We do not know the answer. If this process indeed corresponds to a non-contractible loop in $\mathfrak{m}(X,X_{+})$, it may have nontrivial physical consequences in contexts with topological defects. Recall that references state on annulus $X_{+}$ may come from a system with a defect line that passes through $X_{+}$ as discussed in Section 5.
In summary, the problem we try to formulate in this section is meant to be a pure topology problem. This problem is carefully separated from the study of quantum states on such regions. Nonetheless, we believe this problem is a crucial mathematical problem towards a clear understanding of the automorphism of information convex sets. We wonder if the answers to these questions (or ones that are close analog) have appeared in math literature. The question on $\pi_{0}(\mathfrak{M}(\omega,N))$ is essentially the classification of immersion up to regular homotopy. Some related results can be looked up [^18] Even in this case, we do not always find the results we want. We are most interested in manifolds with boundaries (such as punctured manifolds) immersed in a manifold with the same dimension. Sometimes, a math result we spent time searching for came out fresh [40]. in existing math literature, e.g., [38, 39]. It is unclear to us if the fundamental group $\pi_{1}(\mathfrak{m}(\Omega,N))$, e.g., for examples listed above, has been studied in math literature. (That is why we only provided guesses to most answers!) In any case, entanglement bootstrap provides (adds) a motivation for the study of such mathematical objects.

### v4

@@ 5.2 Constraints of transportation
As a corollary, $|{\mathcal{C}}|=|{\mathcal{C}}_{8}|$. This is because transporting anyon along a figure-8 immersed in a ball or sphere cannot change the anyon type.
Here is a sketch of the key idea of the proof of Prop. 12. The trick is to attach the test annulus to another annulus by a strip. After transporting the test annulus, the strip will be stretched along loop ${\mathcal{L}}$. After gluing the test annulus with the other annulus (by merging with another annulus), we get a punctured torus ${\mathcal{W}}_{\mathcal{L}}$. This process is illustrated in Fig. 14. The main part of the proof is to analyze $\Sigma({\mathcal{W}}_{\mathcal{L}})$, or more conveniently $\Sigma(T^{2})$, where the torus $T^{2}$ has a reference state obtained by an analog of “vacuum block completion” (see Example 4.27 of [17]).
@@
Consider the process illustrated in Fig. 14. We start with a disk away from the defect points (black dots) and remove two small balls from it. Let the resulting 2-hole disk be $Y$. Consider a state labeled by $a,1,a$ on the three boundaries $\rho^{a1a}_{Y}\in\Sigma(Y)$. It must be an extreme point because the corresponding fusion space is one-dimensional.
@@
Figure 14: Illustrated is the construction of an extreme point in $\Sigma_{\hat{1}}({\mathcal{W}}_{\mathcal{L}})$, which is useful in the proof of Prop. 12.
If $a\in{\mathcal{I}}({\mathcal{L}})$, that is, $a$ is invariant under the transportation along ${\mathcal{L}}$, we do step 3 of Fig. 14, which glues annulus $X$ with the state on $Y_{\mathcal{L}}$ to a punctured torus ${\mathcal{W}}_{\mathcal{L}}$. In terms of quantum states, we merge two quantum states, which are extreme points of $\Sigma(X)$ and $\Sigma(Y_{\mathcal{L}})$ (thickening a certain region before merging if needed), with the superselection sectors labeled in Fig. 14(c). Here, $a\in{\mathcal{I}}({\mathcal{L}})$ is required so that the two states match on their overlap regions. For each valid choice of $a$, we get an extreme point of $\Sigma_{\hat{1}}({\mathcal{W}}_{\mathcal{L}})$. Note that $1\in{\mathcal{I}}({\mathcal{L}})$, and therefore $\Sigma_{\hat{1}}({\mathcal{W}}_{\mathcal{L}})$ is nonempty. Note that the punctured torus ${\mathcal{W}}_{\mathcal{L}}$ contains annuli $X_{\mathcal{L}}$ and an embedded test annulus $X$.
@@
where $\delta_{a,b}$ comes from the matching of the sector on one shared boundary of the two annuli, and $\delta_{a,\varphi_{\mathcal{L}}(b)}$ comes from the other shared boundary. The appearance of $\varphi_{\mathcal{L}}$ is due to the transportation property.
Similarly, we cut the torus along another direction into two annuli, $T^{2}=X_{\mathcal{L}}\bar{X}_{\mathcal{L}}$. Here, $X_{\mathcal{L}}$ is the immersed annulus that thickens ${\mathcal{L}}$, a notation explained in Fig. 12. Not only are $X_{\mathcal{L}}$ and $\bar{X}_{\mathcal{L}}$ both sectorizable, but they must also have isomorphic information convex sets $\Sigma(X_{\mathcal{L}})\cong\Sigma(\bar{X}_{\mathcal{L}})$. This is because $X_{\mathcal{L}}$ deforms to $\bar{X}_{\mathcal{L}}$ smoothly on $T^{2}$. So they both characterize ${\mathcal{C}}_{\mathcal{L}}$. From a computation parallel to Eq. (34), we see
@@
The annulus $X_{\mathcal{L}}$ can be viewed as a disk with a puncture. If the state on $X_{\mathcal{L}}$ is an extreme point, we can use the completion trick (Lemma 4.4 of [17]) to heal a puncture of $X_{\mathcal{L}}$ to obtain a state on the disk. The resulting state satisfies the axioms on the entire disk if the extreme point is associated with an Abelian sector ($\mu\in{\mathcal{C}}_{\mathcal{L}}$ with $d_{\mu}=1$). Let this state be $\tilde{\sigma}$. It is known that transporting a test annulus on a disk cannot permute anyon types, supposing that the disk has a reference state. (This is by Lemma 4.3 of [14]). As the loop $X_{\mathcal{L}}$ is identified as part of the disk for which we have a valid reference state $\tilde{\sigma}$, the transportation of the test annulus on $X_{\mathcal{L}}$ cannot permute anyons. Thus, the transportation along ${\mathcal{L}}$ is parallel. ∎
@@ 6 Discussion
In this work, we started by asking a simple question: Is there an Abelian superselection sector detected by the immersed figure-8 annulus?” The question is relevant to gapped topologically ordered systems in 2D, which support anyons and is formulated in the framework of entanglement bootstrap. We gave an affirmative answer to this question for Abelian anyon theories, but leave the general case as a conjecture (Conjecture 8). We explained the fact that if there is an Abelian sector on figure-8, a certain strong isomorphism (Conjecture 9) of information convex sets must be true: if two homeomorphic regions $\Omega$ and $\Omega^{\prime}$ are immersed in the sphere or a disk, the information convex sets $\Sigma(\Omega)$ and $\Sigma(\Omega^{\prime})$ must be isomorphic. Importantly, it does not matter if $\Omega$ can be smoothly deformed to $\Omega^{\prime}$ on the sphere (or disk), a situation where the established isomorphism theorem does not apply.
@@
Suppose the existence of an Abelian sector on the figure-8 (Conjecture 8) can be proved. Strong isomorphism will be a fact. While strong isomorphism is powerful, it seems to imply that we cannot learn much by looking at a topological space immersed differently. Let me argue this is far from the case. If strong isomorphism holds, a couple of intriguing questions can be pursued. In particular, it means tunneling processes (Section 3.1 and Fig. 7) can generate explicit isomorphisms not allowed by smooth deformation. Thus, automorphisms of information convex sets will not only come from topologically nontrivial ways to smoothly deform an immersed region back to itself but also from ways that utilize tunneling in intermediate steps. We expect these diverse topological classes of automorphisms to be informative. A natural question is: “Can we use these automorphisms (dancing of quantum states) to extract all the data of the modular tensor category underlying an anyon model?” [^14] An optimism to this possibility was found independent by Kyle Kawagoe, who attended my chalk talk at math department Ohio State University. If the answer is affirmative, it will manifestly use a single quantum state, and this will be complementary to the “microscopic approach” [40]. Can the mathematical structure coming out of immersion be comparable to the mapping class group? Analogous questions in higher dimensions can be asked in parallel and are more open.
A stronger conjecture we postponed until now is the existence of a canonical (Abelian) vacuum state on the immersed figure-8 annulus. [^15] In some classes of solvable models, such as the quantum double, indeed, such a unique Abelian vacuum on the figure-$8$ can be defined with a technique that depends on the special model. The challenge lies in the general definition. A toy version of the challenge already appears in the chiral semion chiral topological order (assuming the errors of the axioms do not affect the argument). If this is true, the process in Fig. 18 of the appendix may well be on the right track in extracting the full set of anyon topological spins. This is a meaningful problem that deserves an effort elsewhere. Proving the existence of a canonical vacuum on figure-8 remains an open problem, even for Abelian anyon theories, in which context we proved Conjecture 8.
@@
Finally, we ask whether considering quantum states on regions immersed in a physical system can be useful in other physical contexts. Immersed regions utilize the Hilbert space more efficiently by using local pieces more than once, and they can be made larger and topologically (or geometrically) more interesting than the original physical system available to us. Smooth deformation of immersed regions, a property natural in systems with entanglement area law, will be absent in broader contexts such as gapless systems. Nevertheless, many other benefits of immersion may persist.
@@ Appendix A Equivalent definitions of quantum dimension
Consider the context of $n$-dimensional entanglement bootstrap. We derive an equivalence relation on the definition of quantum dimensions (Lemma 15) for immersed sphere shells in $n$-dimensions, where $n\geq 2$ and is arbitrary. It applies to immersed annuli as a special case ($n=2$).
Because the setup is general, we introduce some diagrammatic notation for the partitions. Let $Z=S^{n-1}\times\mathbb{I}$ be a sphere shell in $n$ space dimensions, where $S^{n-1}$ is a $(n-1)$-sphere and $\mathbb{I}$ is an interval. In Fig. 15, the horizontal direction indicates the direction of the interval $\mathbb{I}$. Therefore, each vertical strip that connects the top and the bottom represents a sphere shell. If a strip is cut in half (top and bottom), each piece is a half-sphere shell (topologically a ball). Note that $Z$ is sectorizable, and thus $\Sigma(Z)$ is a simplex. We denote the superselection sectors associated with the extreme points of $\Sigma(Z)$ as ${\mathcal{C}}_{Z}$.
@@
We start with a useful fact. An element $\rho_{Z}$ of $\Sigma(Z)$ is an extreme point if and only if
$\Delta(\partial Z,Z\setminus\partial Z)_{\rho}=0.$ (37)
Here $\partial Z$ is the thickened boundary of $Z$, which is $BD$ in the partition Fig. 15(a). $Z\setminus\partial Z$ is the interior of $Z$, which is $C$ in the same partition. This condition is the extreme point criterion [16, 15].
@@
The following are equivalent definitions of quantum dimension $d_{h}$ of $h\in{\mathcal{C}}_{Z}$, where $Z$ is an immersed sphere shell in $n$-dimensional space.
@@
We shall see in the proof that we do not need to deform the annulus. Therefore, Lemma 15 holds not only for sphere shells immersed in a sphere but also is applicable to the context where the sphere shells are immersed in other topological manifolds.
@@
The extreme point $\rho^{h}_{Z}$ of $\Sigma(Z)$, satisfies $\Delta(\partial Z,Z\setminus\partial Z)_{\rho^{h}}=0$, by the “extreme point criterion”. It implies
@@ Appendix B Automorphism of information convex sets: tools and examples
Automorphism (i.e., self-isomorphism) of information convex sets plays an important role in the entanglement bootstrap. The simplest application is the definition of antiparticles, but its implication is much broader. In this appendix, we talk about two things related to automorphisms of information convex sets.
@@ B.1 Math related to immersion and deformation classes
Here, $\coprod$ means disjoint union (of topological spaces), and $\mathfrak{m}(\Omega,N)$, for any $\Omega\looparrowright N$, is the simplicial 2-complex such that
@@
One may ask if it is meaningful to add higher dimensional $k$-simplexes ($k\geq 3$) instead of stopping at $k=2$. The relevance to physics we can think of so far (as we explain shortly) depends on $\pi_{0}$ and $\pi_{1}$. Adding higher-dimensional simplexes will not change them. Nonetheless, it remains an interesting question if higher homotopy (or homology) groups of $\mathfrak{M}$ can provide further information. [^17] We thank John McGreevy for this question.
@@ B.1.1 Physical relevance of immersion complex through the homotopy groups
It is easy to explain why the number of connected components $\pi_{0}$ of $\mathfrak{M}(\omega,N)$ is of physical relevance. Consider the entanglement bootstrap setup where $N$ is the manifold on which we define our reference state $\sigma_{N}$. The $n$-dimensional version of axioms A0 and A1 are satisfied everywhere on $N$. Then $\pi_{0}(\mathfrak{M}(\omega,N))$, i.e., the number of connected components of $\mathfrak{M}(\omega,N)$, is the number of inequivalent immersions of the topological space $\omega$ in $N$. For instance, $\pi_{0}(\mathfrak{M}(\textrm{annulus},S^{2}))$ has two elements. [^18] In general, $\pi_{0}(\mathfrak{M}(\omega,S^{2}))$ refers to the set of connected components of $\mathfrak{M}(\omega,S^{2})$, and it does not automatically form a group. In the special case that $\omega$ is the annulus, however, there are ways to view the two components as a group $\mathbb{Z}_{2}$. The rest of the paper is independent of whether to make such an assignment. We discussed this in the main text without using this notation. In broader contexts, $\pi_{0}(\mathfrak{M})$ is the set of regular homotopy classes of immersed regions, which has been a topic of study in topology; see, e.g., [41, 42, 30]. Within each connected component, $\mathfrak{m}(\Omega,N)$, $\Omega\in\mathcal{R}(\omega,N)$, we can smoothly deform the region and the isomorphism theorem implies that such deformation preserves the information convex set. Configurations in two connected components are harder to relate, although tunneling (Section 3.1) provides one way to generate isomorphisms between such configurations (when a certain Abelian sector exists).
@@ B.1.2 Examples
where $X(j)$ for $j\geq 0$ represents an immersed annulus with turning number $j$. In particular, we can let $X(0)=X_{8}$ be the immersed “figure-8” and let $X(1)=X$ be the embedded annulus. We expect $\mathfrak{m}(X,B^{2})$ to be simply connected and $\pi_{1}(\mathfrak{m}(X,B^{2}))$ is trivial. [^19] If we deform $X$ back to itself, keeping the intermediate configurations embedded, the deformation must correspond to a loop in $\mathfrak{m}(X,B^{2})$ that can shrink to a point. The argument is precisely that used to prove Lemma 4.3 of [14]. However, for the deformations allowing immersed intermediate configurations, the argument does not apply. What is $\pi_{1}(\mathfrak{m}(X_{8},B^{2}))$? From the way to turn the figure-8 inside out by a $180^{\circ}$ rotation showing in Eq. (12), we see that $\pi_{1}(\mathfrak{m}(X_{8},B^{2}))$ has a quotient group $\mathbb{Z}_{2}$. However, from Ref. [43], we anticipate that $\pi_{1}(\mathfrak{m}(X_{8},B^{2}))=\mathbb{Z}$ [^20] I thank John McGreevy, Joel Hass, and Ryan Budney for the reference.
@@
where $X$ and $X_{8}$ are the only two classes of immersed annuli in the sphere (as discussed in Table 1). Thus $\mathfrak{M}(\textrm{annulus},S^{2})$ has two connected components. What are the fundamental groups $\pi_{1}$ of the connected components? We anticipate that $\pi_{1}(\mathfrak{m}(X,S^{2}))$ has a quotient group $\mathbb{Z}_{2}$. The related observation is that $X$ can turn inside out on the sphere and the fact that such a process can turn anyon $a\in{\mathcal{C}}$ into anti-anyon $\bar{a}\in{\mathcal{C}}$, which is, in general, a different sector. (For a related discussion, see Appendix G of [14].) It is worth understanding if $\pi_{1}(\mathfrak{m}(X,S^{2}))$ itself is $\mathbb{Z}_{2}$. Similiarly, for the figure-8 annulus $X_{8}$, it is worth understanding if $\pi_{1}(\mathfrak{m}(X_{8},S^{2}))$ is $\mathbb{Z}_{2}$; if this is true, the definition of anti-sector on $X_{8}$, considered in Eq. (12) is justified.
@@
where $Z$, on the right-hand side, represents the embedded sphere shell in the 3-dimensional ball $B^{3}$. Eq. (50) follows from the fact that there is a unique way to immerse a sphere $S^{2}$ in $B^{3}$ (or $R^{3}$), up to regular homotopy; this is an influential early result by Smale [41], which is also the topic of video [31]. We are wondering if $\pi_{1}(\mathfrak{m}(Z,B^{3}))=\mathbb{Z}_{2}$; if this is the case, all methods of turning a sphere inside out must be topologically equivalent. It is further interesting to consider
@@
where $G_{g}$ is an embedded genus-$g$ handlebody and $g$ is a positive integer. As is indicated in Eq. (51), $\mathfrak{M}$ is connected in this example. Genus-$g$ handlebodies detect the “graph excitations” studied in [16, 17], where for $g=1$ these graph excitations are flux loops. In general, $\pi_{1}(\mathfrak{m}(G_{g},B^{3}))$ can have nontrivial structure even though $G_{g}$ are sectorizable regions.
Figure 17: Illustrated is a deformation process, which corresponds to an oriented loop in $\mathfrak{m}(X,X_{+})$. Here, $X$ is an annulus (thick black loop in the top figure) embedded in a wider annulus $X_{+}$ (gray).
@@
Note that, $\pi_{1}(\mathfrak{m}(X,X_{+}))$ can be nontrivial even if $\mathfrak{m}(X,B^{2})$ is trivial. Intuitively, in the process described in Fig. 17, a pair of twists is created locally and then transported around the annulus before they annihilate again. (The process in Fig. 17 is inspired by Fig. 8 of [42] as well as Fig. 7 of [20].) Can this process represent a nontrivial element of $\pi_{1}(\mathfrak{m}(X,X_{+}))$? We do not know the answer. If this process indeed corresponds to a non-contractible loop in $\mathfrak{m}(X,X_{+})$, it may have nontrivial physical consequences in contexts with topological defects. Recall that references state on the annulus $X_{+}$ may come from a system with a defect line that passes through $X_{+}$ as discussed in Section 5.
In summary, the problem we try to formulate in this section is meant to be a pure topology problem. This problem is carefully separated from the study of quantum states on such regions. Nonetheless, we believe this problem is a crucial mathematical problem that will help us gain a clear understanding of the automorphism of information convex sets. We wonder if the existing tools in math literature are enough to produce answers to these questions. The question on $\pi_{0}(\mathfrak{M}(\omega,N))$ is essentially the classification of immersion up to regular homotopy. Some related results can be looked up [^21] Even in this case, we do not always find the results we want. We are most interested in manifolds with boundaries (such as punctured manifolds) immersed in a manifold with the same dimension. Sometimes, a math result we spent time searching for came out fresh [44]. in existing math papers e.g., [41, 42, 43]. To our knowledge, the fundamental group $\pi_{1}(\mathfrak{m}(\Omega,N))$, e.g., for examples listed above, has not been studied in the math literature in any systematic way. (That is why we only provided guesses to most answers!) In any case, entanglement bootstrap provides (adds) a motivation for the study of such mathematical objects.